OpenStudy (anonymous):

What equation represents the line passing through the points (2,-2) and (-5,12) ?

OpenStudy (callisto):

Use this: \[\frac{y-y_1}{x-x_1} = \frac{y_2-y_1}{x_2-x_1}\]where \((x_1, y_1)\) and \((x_2, y_2)\) are the coordinates of the points lying on the line.

OpenStudy (anonymous):

Solo 14 over 3...?

OpenStudy (anonymous):

So*

OpenStudy (anonymous):

The answer choices are 2x-y=2 2x+y=2 X+2y=2 X-2y=2

OpenStudy (callisto):

Not really, you have to substitute the values into that form and simplify it. Let say, \((x_1, y_1)\) = (2,-2) and \((x_2, y_2)\) = (-5,12) Sub. the values into it. \[\frac{y-(-2)}{x-2} = \frac{12-(-2)}{(-5)-2}\]^Simplify and rearrange the terms.

OpenStudy (anonymous):

Ya that's 14/3

OpenStudy (anonymous):

@Callisto

OpenStudy (callisto):

No.... the denominator is -5 - 2 and it's not 3...

OpenStudy (anonymous):

Oh -7

OpenStudy (anonymous):

@Callisto

OpenStudy (callisto):

Yup. For the denominator... So, you got \[\frac{y-(-2)}{x-2} = \frac{14}{-7}\]Can you simplify it?

OpenStudy (anonymous):

Well I know 14/-7 simplifies to -2 @Callisto

OpenStudy (callisto):

Yes :) What is y-(-2)?

OpenStudy (anonymous):

Idk cause idk what y is :(

OpenStudy (anonymous):

@Callisto

OpenStudy (callisto):

we.. just simplify it... y-(-2) = y+2

OpenStudy (anonymous):

Oh ok..

OpenStudy (anonymous):

Ok so y+2/x-2 = -2 @Callisto

OpenStudy (callisto):

Yes. Multiply x-2 to both sides, what do you get?

OpenStudy (anonymous):

-2x+4 on the right side @Callisto

OpenStudy (anonymous):

Idk bout the left

OpenStudy (callisto):

For the left, the denominator of the fraction will be cancelled. So, you get y+2 = -2x+4 Rearrange the terms and get your answer :)

OpenStudy (anonymous):

Ok got it. What's the answer I wanna make sure I did it right @Callisto

OpenStudy (callisto):

What have you got?

OpenStudy (anonymous):

2x-y= 2

OpenStudy (anonymous):

Or 2x+y=2 idk

OpenStudy (anonymous):

@Callisto

OpenStudy (callisto):

Which one makes more sense to you??

OpenStudy (anonymous):

The first one ??

OpenStudy (callisto):

Why???

OpenStudy (anonymous):

Cause u have to subtract the y on the other side

OpenStudy (anonymous):

Is it right?

OpenStudy (callisto):

Not really... y+2 = -2x+4 y+2 -y= -2x+4 -y 2 = -2x - y + 4

OpenStudy (anonymous):

That's not even one of the answer choices

OpenStudy (callisto):

You need some more work... Btw, I was showing you ''Cause u have to subtract the y on the other side''

OpenStudy (anonymous):

Ok.... I'm confused lol.. Never mind then.

OpenStudy (callisto):

The best way is to add 2x to both sides instead of subtracting y from both sides. y+2 = -2x+4 y+2 + 2x = -2x + 4 + 2x 2x + y + 2 = 4 <- simplify it.

OpenStudy (anonymous):

Y= 2x+6 ...?

OpenStudy (anonymous):

Or wait its y= 2x+2!!!

OpenStudy (anonymous):

Oh please just tell me the answer ?! :( @Callisto

OpenStudy (callisto):

We are not allowed to tell answers.

OpenStudy (anonymous):

Ok never mind.

OpenStudy (callisto):

When you get to this step 2x + y + 2 = 4 Subtract 2 from both sides. You can probably get your answer then.

OpenStudy (anonymous):

2x+y=2??!!???!

OpenStudy (anonymous):

Got it thank you.